A perfectly competitive industry consist of 300 firms with identical cost structures. Their respective market demand and market supply functions are given as Qd= 3000 - 60p and Qs= 500 + 40p.\na) what are the profit maximizing price and output for each individual firm.\nb) assume that firm's in the long run competitive equilibrium, determine each firm's TR, TC and profit.


1
Expert's answer
2021-04-02T11:21:05-0400

na)

"3000-60p=500+40p"

"100p=2500"

"p=25"

"Q=\frac{1500}{300}=5"

nb)


"TR=25\times50=1250"

"\pi=TR-TC"